COMPARISON OF LOW TEMPERATURE PRESSURE BROADENING OF NH_{3} BY HE, NORMAL- H_{2} AND PARA- H_{2}

\underline{D. R. WILLEY}, R. E. TIMLIN, JR., J. M. MERLIN, M. M. SOWA and D. M. WESOLEK, Department of Physics, Allegheny College, Meadville, PA 16335.

We have completed low temperature pressure broadening studies of NH_{3} inversions transitions for broadening by He and molecular hydrogen. Pressure broadening of the (J, K) = (1, 1), (2, 2) and (3, 3) inversion transitions of NH_{3} by He, normal- H_{2} and para- H_{2} has been observed from 10 to 40 K. We find that the He cross sections are the smallest followed by the para- H_{2} cross sections which are two to five times larger with the normal- H_{2} cross sections being two to ten times larger than for He. In addition, we have used ab initio potential surfaces to calculate He and para- H_{2} cross sections for comparison with our experimental data.